# LEP3: A High-Luminosity $e^{+}e^{-}$ Higgs & Electroweak Factory in the LHC Tunnel

### v1

@@
April 1, 2025
C. Anastopoulos1, R. Assmann16, A. Ball2, O. Bruning3, O. Buchmueller4, T. Camporesi5,15, P. Collier3, J Dainton6,14, G. Davies4, J.R. Ellis3,7, B. Goddard3, L. Gouskos8, G. Hall4, M. Klute9, M. Koratzinos10, G. Landsberg8, K. Long4, L. Malgeri3, F. Maltoni11,17, F. Moortgat3, C. Mariotti12, S. Myers3, J.A. Osborne3, M. Pierini3, D.R. Tovey1, D. Treille3, T.S. Virdee4, N. Wardle4, M. Zanetti13
@@ 1 Introduction
To maintain a high instantaneous luminosity $(1.5-2.5\times 10^{34}cm^{-2}s^{-1}$ at $\sqrt{s}=230\,GeV$), separate rings are required for the collider and the full energy booster (accelerator), the latter for top-up injection.
@@
The baseline anticipates two experiments. During the running programme, around $\approx 5\times 10^{5}\,e^{+}e^{-}$ ZH events would be recorded over 6 years at $\sqrt{s}=230\,GeV$. LEP3 would also run at the Z and WW thresholds, recording $\approx 2\times 10^{12}$ Z decays over 5 years at or around $\sqrt{s}=91.2\,GeV$, and $\approx 4\times 10^{7}$ WW events at or around $\sqrt{s}=163\,GeV$ over four years.
@@
Using the FCC-ee cost methodology, the cost of LEP3, including costs related to civil engineering, LHC removal, LEP3 installation and two new experiments is estimated to be around 3.2 BCHF for all stages. The cost of an initial stage running at the Z peak would be $\approx$2.7 BCHF. In the costs presented here we have included the cost of two new experiments at $\approx$0.5 BCHF each with CERN’s contribution amounting to 10% of 1 BCHF. The additional cost to the community would thus be $\approx$0.9 BCHF. If the cost has to be limited, the existing ATLAS and CMS experiments could be re-deployed, suitably modified/upgraded.
We consider that the prospective physics performance of LEP3 is fully competitive with other proposed alternatives to the preferred FCC option at a lower cost. However, LEP3 is not competitive with FCC-ee.
@@ 2 The Main Stages of LEP3
The LEP3 physics programme would have three phases,
  • near or on the Z peak (91.2 $GeV$), over 5 years,
  • near the WW threshold (163 $GeV$), over 4 years and
  • near the ZH threshold (at 230 $GeV$), over 6 years.
The broad aim is to conduct precision electroweak physics involving the W and Z bosons as well as precision Higgs boson physics. The order proposed above corresponds to increasing the energy monotonically, but another ordering could also be considered (e.g., 6 years at ZH, 5 years at and around the Z-peak, and 4 years at the WW threshold). Running at increasing energy steps has the advantage of spreading out the required expenditure. The duty cycle assumed is the same as for FCC-ee, i.e., 185 days @ 75% efficiency giving an effective running time of $1.2\times 10^{7}$ seconds per year.
@@ 3 Technical Considerations
Luminosity and CoM Energy for ZH Operation: The maximum CoM energy has been chosen to be 230 $GeV$. Although the total Higgs boson cross section peaks at a CoM energy of 250 $GeV$, the luminosity in a circular collider, at a fixed synchrotron radiation power loss, drops rapidly with beam energy, moving the peak rate of Higgs bosons lower (to 235 $GeV$ in the case of LEP3/FCC-ee). The luminosity of such a collider would need a design of a low-emittance lattice and detailed simulations. In their absence, we give a range of the luminosity expected at $\sqrt{s}=230\,GeV$ that is between $1.5-2.5\times 10^{34}cm^{-2}s^{-1}/IP$ for two IPs. Simple scaling of luminosity from FCC would give a factor  3.5 lower luminosity. We have also performed a more detailed but preliminary analysis that can be found in the accompanying document [3]. To calculate the number of events for the physics analyses we have used a luminosity value of $1.8\times 10^{34}cm^{-2}s^{-1}/IP$ at $\sqrt{s}=230\,GeV$, $6.4\times 10^{34}cm^{-2}s^{-1}/IP$ at or around $\sqrt{s}=163\,GeV$, $44\times 10^{34}cm^{-2}s^{-1}/IP$ at or around $\sqrt{s}=91.2\,GeV$. Whereas the rates of Higgs boson production at linear colliders would be similar to that potentially achievable at LEP3 for a single IP, the luminosities are considerably smaller near the Z peak and the WW threshold. Consequently, LEP3 has a much stronger precision electroweak programme, complementing its precision Higgs physics one.
@@
Energy Calibration: It is important for precision electroweak physics to measure the beam energy as precisely as possible for Z and WW running. The contribution of the beam energy uncertainty on the W mass measurement uncertainty at LEP, compared with that of the Z mass mesurement, was a factor 10 higher (25 $MeV$) as no resonant depolarisation measurements were possible at the WW energies (little polarisation was observed). For LEP3, we can use the FCC-ee approach for Z running, based on the knowledge that LEP routinely achieved adequate polarisation levels at the Z. For WW running where no polarization was observed at LEP, precise simulations will be needed to see if adequate levels of polarization can be achieved at energies close to or at the WW threshold. The benefits available with modern technologies, that were not available for LEP1 and LEP2, need to be investigated in addition. For example, much lower than 5% polarisation levels might be sufficient for resonant depolarisation with today’s technologies. Use of modern diagnostics and electronics will thus likely enable energy calibration at higher energies. Failing this, the aim is to use the resonant depolarisation method at the highest energy possible, and extrapolate to the WW threshold. This needs further study.
Beam Energy Spread: In the absence of beamstrahlung Beam Energy Spread, $\sigma(E_{b})\propto E_{b}^{2}\sqrt{\rho}$ where $E_{b}$ is the beam energy, and $\rho$ is the bending radius. Our estimate is that the total beam energy spread at $E_{b}=m(Z)$ at LEP3 should be similar to that for FCC-ee. The contribution of synchrotron radiation is smaller at the FCC but the beamstrahlung contribution is larger because the instantaneous luminosity is higher.
@@
Crossing Angle and Implications: A “crab-waist” scheme with a significant crossing angle of around 30 $mrad$ (half-angle of 15 $mrad$) is necessary to attain high instantaneous luminosity [4]. Such large crossing angles have implications for the exact location of the interaction point (IP) and for the paths of the beams in the straight sections in the vicinity of the IPs. A scheme similar to that deployed for CEPC [5] can be found for LEP3, in which the IP can be kept at the coordinates of the current IPs in ATLAS and CMS, and limiting the spatial excursion of the beams in the straight sections. Studies are needed to see exactly what would be achievable at LEP3.
@@
Booster Bypass: Consideration has to be given to the booster/accelerator bypass around the experiments, possibly still remaining inside the experiment caverns, e.g. running along the balconies
Injector: Arguments for a dedicated linac injector, rather than using the SPS as injector, are given in the FCC MTR [2]. Therefore, here we pursue only the linac option. The injection energy is affected by non-uniformities in the bending magnets at the lowest magnetic field i.e. at injection energy. Although the injection energy potentially can be lowered by the ratio of the circumferences, i.e., by a factor of 3.4, we assume a higher injection energy of 10 $GeV$. This is to be compared with 20 $GeV$ for FCC-ee. The lower energy allows the linac length to be halved compared to what would be required for FCC-ee, while keeping a similar energy swing as the FCC. We adopt the siting study carried out for FCC-ee to locate the linac injector on the Prevessin site, but the shorter length allows for a more optimal placement minimizing transfer tunnel length. The layout can still be optimized to reduce its length or use existing tunnels.
Synchrotron Power Loss/Turn and RF Requirements: The energy loss/turn for LEP3 operating at 230 $GeV$, with a bending radius like that at LEP2, is 5.4 $GeV$/turn. With a margin, this leads to a requirement of 6 $GV$ of RF to be installed, compared with 2.1 $GV$ for FCC-ee.
@@
Civil Engineering Works: The LEP/LHC tunnel was constructed in the 1980s with an assumed lifetime of some 40 years. To prolong its usability, either for LEP3 or as an injector for the future FCC-hh programme, maintenance will be needed in some sections considered to be fragile. It is estimated that this repair corresponds to a section of a length of 0.6 $km$ in Sector 3-4. Furthermore, the straight sections around the two collision IPs will need widening to accommodate the large crossing angle. The length to be widened corresponds to two times 2$\times$270 $m$, with a maximum diameter of  7 $m$. Allowance has been made for some additional cores from the klystron galleries into the tunnel. It is assumed that the existing LHC shafts can be used to carry out the LEP3 civil engineering works. The civil engineering works are estimated to cost around 165 MCHF.
@@
Radioprotection Issues: Studies have been carried out in the context of HL-LHC that indicate that the activation levels are low inside the LHC tunnel (surface of the tunnel walls and machine elements inside). We do not anticipate any problems in the arcs but further studies would be needed before drawing firm conclusions on how the dismantling work should proceed.
Experiments: Although having two new experiments in the existing ATLAS and CMS caverns may be preferable for many reasons, including a better-adapted physics performance, also use could be made of the two existing LHC general-purpose experiments i.e. ATLAS and CMS suitably modified or upgraded, especially the inner tracking systems. Which option to adopt will require an optimization and considerations concerning the cost, the physics performance and sociology. If this Project proceeds it is probable that a call for experiments would be made and new and/or suitably upgraded existing experiments inevitably would be proposed. A previous study explored the suitability of the CMS experiment for such e+e- operation [6].
The R&D carried out by the ALICE collaboration for its new bendable pixel detector could form a possible basis for LEP3 pixel and tracking detectors. Special luminometers will be required. If existing detectors are used, finer lateral granularity, e.g., for hadronic calorimeters may have to be introduced. The TDRs for the experiments (new or re-deployed) would have to be submitted by the mid-2030s, several years before the start of the shutdown for the transition to LEP3, enabling a head start for construction. The trackers would be designed to be light-weight, with a momentum resolution for 40 $GeV$ charged particles at the $\sim 0.1$% level, and an angular resolution on 40 $GeV$ muons of $\sim 0.1\,mrad$.
If ATLAS is reused, modifications to the endcap toroid magnets and/or liquid argon calorimeter may be required to accommodate the final focus of the collider ring.
Some level of particle identification can be obtained from the precision timing detectors being installed for HL-LHC running. However, unless the possibility of two new detectors is adopted, probably there is no room to introduce significant $\pi$/K separation capability.
We estimate that two new experiments would cost the community some 2$\times$0.5 BCHF. The cost to be borne by CERN is assumed to be 10%, as in the case of FCC-ee costing, ie. 0.1$\times$2$\times$0.5 BCHF = 100 MCHF. This has been included in the cost estimate presented here. In existing detectors, new inner trackers would cost around 100 MCHF per experiment, and we estimate that another 50 MCHF would be needed for all other modifications.
It is assumed that a large fraction of the existing experimental collaborations would remain to exploit data from LEP3. If ATLAS and CMS are re-deployed there would be a considerable advantage in using understood detectors, software and analysis tools, and scientists already experienced in physics analysis of data from these existing experiments.
For Z running the inner detector would be immersed in a solenoid field of 2T in both the ATLAS and CMS experiments. In CMS the solenoid magnetic field could be raised for higher $\sqrt{s}$ running, e.g., 3 Tesla during ZH running. In ATLAS it is expected that the toroid magnets for muon momentum measurement will continue to operate, possibly at a reduced field. The rest of the ATLAS and CMS detectors should continue to function well during the duration of LEP3 programme.
Machine-Detector Interface: There would be superconducting final quadrupoles near the interaction regions that follow the design outlined in the FCC MTR [2]. Starting at a distance of 2.2 $m$ from the interaction point, a shielding solenoid surrounds the quadrupole, and a compensating solenoid is located close to the collision point. A luminometer is integrated. A stay-clear cone is defined, above 100 $mrad$, from the IP along the z axis. This corresponds to a line with pseudo rapidity of $|\eta|\sim 3$. Thus the forward or very forward calorimeters will not be required.
@@
Cost Estimate: The capital cost of LEP3 has been evaluated employing the methodology used to estimate the cost of FCC-ee, with appropriate scaling (e.g., by the numbers of components) of the costs of individual FCC-ee items. The error in this estimate would be similar but no smaller than that for the FCC-ee. Included in this estimate are the costs of the civil engineering works described above, the cost of careful removal and storage of LHC components to allow possible future re-use, and the installation of the LEP3 machine. The latter costs have been estimated by discussions with those involved in the installation of the original LEP accelerator. The total cost to CERN of the LEP3 project, with two new experiments, is estimated to be 3.2 BCHF (see Table 1). ”Cost to PP” column gives the additional cost to the particle physics community (0.9 BCHF). Redeploying existing experiments leads to a cost of $\sim 3.15$ BCHF to CERN and an additional cost of 270 MCHF to the particle physics community. The cost is dominated by the cost of the RF system ($\sim 0.75$ BCHF). The next costliest items are the quadrupole/sextupole magnets and the injector chain. Operation with progressively increasing CoM energy, i.e., starting with the Z-peak running, could lead to a lowering of the initial cost by some 0.5 BCHF. The FCC costing study uses a classification matrix published by the Association for the Advancement of Cost Engineering (AACE).
@@
Table 1: The cost estimate for LEP3 using the FCC-ee costing methodology including two new experiments (3185 MCHF). Also listed is the additional cost (900 MCHF) to the wordwide particle physics community “Cost to PP”. The same costs for two existing experiments amount to 3135 MCHF and 270 MCHF respectively.
@@
Schedule:The LEP3 accelerator would be installed after the end of the HL-LHC programme. We estimate that five years will be needed for dismantling the LHC, carrying out civil engineering works and installation and commissioning of the LEP3 accelerator. Assuming the current HL-LHC schedule LEP3 could be operational in the mid-2040s.
@@
Sustainability: We assume that all building permissions/permits for LEP3 should either be in place or be easy to obtain. Also, the amount of preparatory work and the environmental impact should be minimal. However, consideration will have to be given, and allowance made, for regulations that may have changed since the time of the construction of LEP. As relatively little construction is required, it is expected that the carbon cost would be a fraction of any new build. Few other significant environmental impacts are anticipated. Nevertheless, benefit would be drawn from any relevant studies in these areas carried out for other proposed projects such as FCC-ee or CLIC.
Power Consumption: We estimate that the total power consumption at the highest LEP3 energy will be about 250 $MW$, like that of HL-LHC. Usage of high-temperature superconducting arc quadrupoles and sextupoles (HTS) magnets would help in minimising this figure.

### v2

@@
Updated on May 27, 2025
C. Anastopoulos1, R. Assmann16, A.H. Ball2, O. Bruning3, O. Buchmueller4, T. Camporesi5,15, P. Collier3, J Dainton6,14, G. Davies4, J.R. Ellis3,7, B. Goddard3, L. Gouskos8, G. Hall4, M. Klute9, M. Koratzinos10, G. Landsberg8, K. Long4, L. Malgeri3, F. Maltoni11,17, F. Moortgat3, C. Mariotti12, S. Myers3, J.A. Osborne3, M. Pierini3, D.R. Tovey1, D. Treille3, T.S. Virdee4, N. Wardle4, M. Zanetti13
@@ 1 Introduction
To maintain a high instantaneous luminosity separate rings are required for the collider and the full energy booster (accelerator), the latter for top-up injection.
@@
The baseline anticipates two experiments. During the running programme, around $\approx 4\times 10^{5}\,e^{+}e^{-}$ ZH events would be recorded over six years at $\sqrt{s}=230\,GeV$. LEP3 would also run at the Z and WW thresholds, recording $\approx 2\times 10^{12}$ Z decays over six years at or around $\sqrt{s}=91.2\,GeV$, and $\approx 4\times 10^{7}$ WW events at or around $\sqrt{s}=160\,GeV$ over four years.
@@
Using the FCC-ee cost methodology, the cost to CERN of LEP3, including costs related to civil engineering, LHC removal, LEP3 installation and two new experiments is estimated to be around 3.9 BCHF for all stages. Some 0.9 BCHF from non-CERN contributions in not included in the above number (see Table 3). In the costs presented here we have included the cost of two new experiments at $\approx$0.5 BCHF each with CERN’s contribution amounting to 10% of 1 BCHF. If the cost to teh community has to be limited, the existing ATLAS and CMS experiments could be re-deployed, suitably modified/upgraded.
We consider that the prospective physics performance of LEP3 is fully competitive with other proposed alternatives, to the preferred FCC option, at a lower cost. However, LEP3 is not competitive with FCC-ee.
@@ 2 The Main Stages of LEP3
The LEP3 physics programme would have three phases, that should include a couple of long shutdowns,
  • near or on the Z peak (91.2 $GeV$), over six years,
  • near the WW threshold (163 $GeV$), over four years and
  • near the ZH threshold (at 230 $GeV$), over six years.
The broad aim is to conduct precision electroweak physics involving the W and Z bosons as well as precision Higgs boson physics. The order proposed above corresponds to increasing the energy monotonically, but another ordering could also be considered (e.g., six years at ZH, six years at and around the Z-peak, and four years at the WW threshold). Running at increasing energy steps has the advantage of spreading out the required expenditure. The duty cycle assumed is the same as for FCC-ee, i.e., 185 days @ 75% efficiency giving an effective running time of $1.2\times 10^{7}$ seconds per year.
@@ 3 Technical Considerations
Luminosity and CoM Energy for ZH Operation: The maximum CoM energy has been chosen to be 230 $GeV$. Although the total Higgs boson cross section peaks at a CoM energy of 250 $GeV$, the luminosity in a circular collider, at a fixed synchrotron radiation power loss, drops rapidly with beam energy, moving the peak rate of Higgs bosons lower (to 235 $GeV$ in the case of LEP3/FCC-ee).
In this update of our submission  [3], we base our parameters on a study made by K Oide and D. Shatilov in 2017 [4] using a preliminary lattice for LEP3, inspired by the FCC-ee design. The Oide/Shatilov study was carried out for four IPs and at $\sqrt{s}=240\,GeV$. Their estimate of the instantaneous luminosity/IP was $1.1\times 10^{34}cm^{-2}s^{-1}$ at $\sqrt{s}=240\,GeV$ and $52\times 10^{34}cm^{-2}s^{-1}/IP$ around the Z-pole. The LEP3 baseline anticipates two IPs and ZH running at $\sqrt{s}=230\,GeV$, leading to an instantaneous luminosity/IP of $1.5\times 10^{34}cm^{-2}s^{-1}$ for ZH running. This increase, from 1.1 to 1.5×$\times 10^{34}cm^{-2}s^{-1}/IP$, is due to two factors of 1.15, corresponding to two experiments and a lowered $\sqrt{s}$ energy, respectively. The Oide/Shatilov study used very strong normal quadrupoles and sextupoles which would lead to a very high power consumption and would need a wide range of hardware site-modifications. For these reasons we propose the use of nested superconducting quadrupoles/sextupoles. This possibility is an option in the FCC-ee design [2][5]. This would result in not only a significantly lower power consumption, but also an increase in instantaneous luminosity as the magnetic bending radius could be 7 % higher (2958m instead of 2755m), reducing the RF required as well as the increasing the luminosity by 7%. We have re-evaluated the LEP3 parameters in the light of this and the Oide/Shatilov study. The parameters from our study are shown in Table 1. We estimate that the maximum RF voltage to be installed will be around 6 GV and the instantaneous luminosity/IP at $\sqrt{s}=230\,GeV$ to be $1.6\times 10^{34}cm^{-2}s^{-1}$. For the Z running, we have changed parameters to respect two aspects: the spin modulation index, a figure of merit for spin depolarization measurements, and the threshold of x-z instability (given by $Q_{s}/\xi_{x}$ ). For the former, the problem was that the momentum compaction factor is much smaller compared to LEP (and similar to FCC). To compensate for that, we have increased $Q_{s}$ by increasing RF voltage. Then, to keep a reasonable beam-beam vertical tune shift, we increased the vertical emittance and decreased the number of bunches. This in turn helped in increasing the bunch length (which is good for x-z instability, resistive wall power, etc.)
Table 1: Accelerator parameters for LEP3
To calculate the number of events for the physics analyses we have used a luminosity value of $1.6\times 10^{34}cm^{-2}s^{-1}/IP$ at $\sqrt{s}=230\,GeV$, $6.2\times 10^{34}cm^{-2}s^{-1}/IP$ at or around $\sqrt{s}=160\,GeV$, $40\times 10^{34}cm^{-2}s^{-1}/IP$ at or around $\sqrt{s}=91.2\,GeV$. Hence, for two experiments, around $4.2\times 10^{5}e^{+}e^{-}\rightarrow ZH$ events would be recorded over six years at $\sqrt{s}=230\,GeV,1.7\times 10^{12}Z$ decays over six years at or around $\sqrt{s}=91.2GeV$, and $3.6\times 10^{7}WW$ events at or around $\sqrt{s}=160\,GeV$ over four years. Whereas the rates of Higgs boson production at linear colliders would be similar to that potentially achievable at LEP3 for a single IP, the luminosities are considerably smaller near the Z peak and the WW threshold. Consequently, LEP3 has a much stronger precision electroweak programme, complementing its precision Higgs physics one.
@@
Energy Calibration: It is important for precision electroweak physics to measure the beam energy as precisely as possible for Z and WW running. The contribution of the beam energy uncertainty on the W mass measurement uncertainty at LEP, compared with that of the Z mass mesurement, was a factor 10 higher (25 $MeV$) as no resonant depolarisation measurements were possible at the WW energies (no polarisation was detected). For LEP3, we can use the FCC-ee approach for Z running, based on the knowledge that LEP routinely achieved adequate polarisation levels at the Z. For WW running where no polarization was detected at LEP, precise simulations will be needed to see if adequate levels of polarization can be achieved at energies close to or at the WW threshold. The benefits available with modern technologies, that were not available for LEP1 and LEP2, need to be investigated in addition. For example, much lower than 5% polarisation levels might be sufficient for resonant depolarisation with today’s technologies. Use of modern diagnostics and electronics will thus likely enable energy calibration at higher energies. Failing this, the aim is to use the resonant depolarisation method at the highest energy possible, and extrapolate to the WW threshold. We also propose to install a large amount of RF voltage for the WW running (3GV compared to an energy loss per turn of 1.5 $GeV$. This needs further study.
@@
Crossing Angle and Implications: A “crab-waist” scheme with a significant crossing angle of around 30 $mrad$ (half-angle of 15 $mrad$) is necessary to attain high instantaneous luminosity [6]. Such crossing angles necessitate special optics for the paths of the beams in the vicinity of the IPs. A scheme similar to that deployed for CEPC [7] can be found for LEP3, in which the IP can be kept at the coordinates of the current IPs in ATLAS and CMS, and limiting the spatial excursion of the beams in the straight sections. We have initiated a study with experts to realise this possibility.
@@
Booster Bypass: Consideration has to be given to the booster/accelerator bypass around the experiments, possibly still remaining inside the experiment caverns, e.g. running along the balconies in the experiment caverns.
Injector: Arguments for a dedicated linac injector, rather than using the SPS as injector, are given in the FCC MTR [2]. Therefore, here we pursue only the linac option. The injection energy is affected by non-uniformities in the bending magnets at the lowest magnetic field i.e. at injection energy. Although the injection energy potentially can be lowered by the ratio of the circumferences, i.e., by a factor of 3.4, we assume a smaller factor, and a higher injection energy of 10 $GeV$. This is to be compared with 20 $GeV$ for FCC-ee. The lower energy allows the linac length to be halved compared to what would be required for FCC-ee, while keeping a similar energy swing as the FCC. We adopt the siting study carried out for FCC-ee to locate the linac injector on the Prevessin site, but the shorter length allows for a more optimal placement minimizing transfer tunnel length. The layout can still be optimized to reduce its length or use existing tunnels.
@@
Civil Engineering Works: The LEP/LHC tunnel was constructed in the 1980s with an assumed lifetime of some 40 years. To prolong its usability, either for LEP3 or as an injector for the future FCC-hh programme, maintenance will be needed in some sections considered to be fragile. It is estimated that this repair corresponds to a section of a length of 0.6 $km$ in Sector 3-4. Furthermore, the straight sections around the two collision IPs will need widening to accommodate the large crossing angle. The length of the tunnel assumed to be widened corresponds to two times 2$\times$270 $m$, with a maximum diameter of  7 $m$. Allowance has been made for some additional cores from the klystron galleries into the tunnel. It is assumed that the existing LHC shafts can be used to carry out the LEP3 civil engineering works. The civil engineering works are estimated to cost around 165 MCHF.
@@
Radioprotection Issues: Studies have been carried out in the context of HL-LHC that indicate that the activation levels inside the LHC tunnel are low (surface of the tunnel walls and machine elements inside). Discussions with CERN’s RP Group suggest that there are no showstoppers but further studies would be needed before drawing firm conclusions and on how the dismantling work should proceed.
Experiments: Although having two new experiments in the existing ATLAS and CMS caverns may be preferable for many reasons, including a better-adapted physics performance, also use could be made of the two existing LHC general-purpose experiments i.e. ATLAS and CMS suitably modified or upgraded, especially their inner tracking systems. Which option to adopt will require an optimization and considerations concerning the cost, the physics performance and sociology. If this Project proceeds it is probable that a call for experiments would be made and new and/or suitably upgraded existing experiments inevitably would be proposed. A previous study explored the suitability of the CMS experiment for such e+e- operation [8]. The R&D carried out by the ALICE collaboration for its new bendable pixel detector could form a possible basis for LEP3 pixel and tracking detectors. Special luminometers will be required. If existing detectors are used, finer lateral granularity, e.g., for hadronic calorimeters may have to be introduced. The TDRs for the experiments (new or re-deployed) would have to be submitted by the mid-2030s, several years before the start of the shutdown for the transition to LEP3, enabling a head start for construction. The trackers would be designed to be light-weight, with a momentum resolution for 40 $GeV$ charged particles at the $\sim 0.1$% level, and an angular resolution on 40 $GeV$ muons of $\sim 0.1\,mrad$. If ATLAS is reused, modifications to the endcap toroid magnets and/or liquid argon calorimeter may be required to accommodate the final focus of the collider ring. Some level of particle identification can be obtained from the precision timing detectors being installed for HL-LHC running. However, unless the possibility of two new detectors is adopted, probably there is no room to introduce significant $\pi$/K separation capability. We estimate that two new experiments would cost the community some 2$\times$0.5 BCHF. The cost to be borne by CERN is assumed to be 10%, as in the case of FCC-ee costing, ie. 0.1$\times$2$\times$0.5 BCHF = 100 MCHF. This has been included in the cost estimate presented here. In existing detectors, new inner trackers would cost around 100 MCHF per experiment, and we estimate that another 50 MCHF would be needed for all other modifications. It is assumed that a large fraction of the existing experimental collaborations would remain to exploit data from LEP3. If ATLAS and CMS are re-deployed there would be a considerable advantage in using understood detectors, software and analysis tools, and scientists already experienced in physics analysis of data from these existing experiments. For Z running the inner detector would be immersed in a solenoid field of 2T in both the ATLAS and CMS experiments. In CMS the solenoid magnetic field could be raised for higher $\sqrt{s}$ running, e.g., 3 Tesla during ZH running. In ATLAS it is expected that the toroid magnets for muon momentum measurement will continue to operate, possibly at a reduced field. The rest of the ATLAS and CMS detectors should continue to function well during the duration of LEP3 programme.
[…]
@@
Cost Estimate: The capital cost of LEP3 has been evaluated employing the methodology used to estimate the cost of FCC-ee, with appropriate scaling (e.g., by the numbers of components) of the costs of individual FCC-ee items. The error in this estimate would be similar to that for the FCC-ee. Included in this estimate are the costs of the civil engineering works described above, the cost of careful removal and storage of LHC components to allow possible future re-use, and the installation of the LEP3 machine. The latter costs have been estimated by discussions with those involved in the installation of the original LEP accelerator. The total cost to CERN of the LEP3 project, with two new experiments, is estimated to be 3.9 BCHF (see Table 3). The cost is dominated by the cost of the RF system ($\sim 1.5$ BCHF). The next costliest items are the quadrupole/sextupole magnets ($\sim 0.4$ BCHF) and the injector chain ($\sim 0.3$ BCHF). The FCC costing study uses a classification matrix published by the Association for the Advancement of Cost Engineering (AACE).
@@
Table 3: The LEP3 cost to CERN is estimated to be 4785 MCHF using the FCC-ee costing methodology including the additional cost of 900 MCHF to the wordwide particle physics community. The same costs for two existing experiments amount to 4085 MCHF including 270 MCHF respectively. These costs are dominated by the cost of the RF system at over 1.5 BCHF
@@
Schedule:The LEP3 accelerator would be installed after the end of the HL-LHC programme. We estimate that five years will be needed for dismantling the LHC, carrying out civil engineering works and installation and commissioning of the LEP3 accelerator. Assuming the current HL-LHC schedule, LEP3 could be operational in the mid-2040s.
R&D Needs: We believe that around 5% of the capital cost of the Project will have to be spent for the remaining R&D and prototyping before proceeding to construction. This would amount to about 200 MCHF. In addition to the R&D needs for the FCC-ee project outlined in submission  [5], two particular areas of R&D affecting LEP3 are called out: R&D on higher power fundamental power couplers, and on nested superconducting quadrupole/sextupole magnets.
@@
Sustainability: We assume that all building permissions/permits for LEP3 should either be in place or be easy to obtain. Also, the amount of preparatory work and the environmental impact should be minimal. As relatively little construction is required, it is expected that the carbon cost would be a fraction of any new build. Few other significant environmental impacts are anticipated. Nevertheless, benefit would be drawn from any relevant studies in these areas carried out for other proposed projects such as FCC-ee or CLIC. Usage of high-temperature superconducting arc quadrupoles and sextupoles (HTS) magnets would help limit the power consumption compared to using warm magnets.
Power Consumption: We estimate that the total power consumption at the highest LEP3 energy will be about 250 $MW$, and about 185 $MW$ at the Z pole.